### v2

# DeepAgent: A General Reasoning Agent with Scalable Toolsets

Abstract.
Large reasoning models have demonstrated strong problem-solving abilities, yet real-world tasks often require external tools and long-horizon interactions. Existing agent frameworks typically follow predefined workflows, which limit autonomous and global task completion. In this paper, we introduce DeepAgent, an end-to-end deep reasoning agent that performs autonomous thinking, tool discovery, and action execution within a single, coherent reasoning process. To address the challenges of long-horizon interactions, particularly the context length explosion from multiple tool calls and the accumulation of interaction history, we introduce an autonomous memory folding mechanism that compresses past interactions into structured episodic, working, and tool memories, reducing error accumulation while preserving critical information. To teach general-purpose tool use efficiently and stably, we develop an end-to-end reinforcement learning strategy, namely ToolPO, that leverages LLM-simulated APIs and applies tool-call advantage attribution to assign fine-grained credit to the tool invocation tokens. Extensive experiments on eight benchmarks, including general tool-use tasks (ToolBench, API-Bank, TMDB, Spotify, ToolHop) and downstream applications (ALFWorld, WebShop, GAIA, HLE), demonstrate that DeepAgent consistently outperforms baselines across both labeled-tool and open-set tool retrieval scenarios. This work takes a step toward more general and capable agents for real‑world applications. The code and demo are available at https://github.com/RUC-NLPIR/DeepAgent.
Large Reasoning Models, Autonomous Agents, Tool Retrieval, Memory Mechanism, Reinforcement Learning
* Work done during internship at Xiaohongshu.
$\dagger$ Corresponding author.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792460ccs: Computing methodologies Planning and scheduling
Figure 1. Overall performance on (a) general tool usage tasks and (b) downstream applications (best score as 100%).
Figure 2. Comparison of agent paradigms: (a) Traditional agents with predefined workflows, (b) Deep Research agents that can autonomously call limited tools, and (c) Our DeepAgent, a fully autonomous reasoning agent that dynamically discovers and invokes helpful tools, all within a continuous agentic reasoning process.
1. Introduction
The rapid advancement of Large Language Models (LLMs) has inspired the development of LLM-powered agents, which have found broad applications in scenarios such as web information seeking, software engineering, and personal assistance (Qu et al., 2025b; Jin et al., 2024). Most existing agents follow predefined workflows (e.g., ReAct (Yao et al., 2022b) and Plan-and-Solve (Wang et al., 2023)) with iterative “Reason-Act-Observe” loops (Figure 2(a)). Although effective in simpler tasks, these approaches suffer from several critical limitations: (1) lack of autonomy in execution steps and overall procedure; (2) inability to dynamically discover tools during task execution; (3) deficiency in fully autonomous management of interactive memory; and (4) insufficient depth and coherence in reasoning about the entire task. These limitations hinder agents from real-world problems, particularly for complex tasks that demand general and multiple tool-use.
Recently, the advent of Large Reasoning Models (LRMs) has demonstrated the capability to solve complex problems in domains like mathematics, programming, and scientific reasoning through a step-by-step “slow thinking” process (Chen et al., 2025a; Li et al., 2025e; Wang et al., 2025c). However, many real-world tasks necessitate the use of external tools for their completion. Recent approaches (e.g., Search-o1 (Li et al., 2025a), DeepResearcher (Zheng et al., 2025), and ToRL (Li et al., 2025d)) integrate tool use into reasoning, but typically rely on a small, fixed tool set such as search, browsing, and coding (Figure 2(b)), limiting their generality.
To address these challenges, we introduce DeepAgent, an end-to-end deep reasoning agent that can complete an entire task by dynamically retrieving and calling tools within a single, coherent agentic reasoning process. As depicted in Figure 2(c), DeepAgent operates by autonomously thinking, searching for tools, and executing actions. This paradigm shifts away from traditional, predefined workflows that rely on predefined tools, task planning, and iterative tool use. Instead, DeepAgent maintains a global perspective on the entire task, unconstrained by the need to deliberate on specific, isolated operations. Tools are not pre-retrieved in advance but are dynamically discovered on an as-needed basis, thereby fully unlocking the autonomous potential of the large reasoning model.
To empower DeepAgent to thoroughly and robustly explore new tools and navigate complex environments during long-horizon interactions, we equip it with memory management capabilities. We introduce an Autonomous Memory Folding strategy that allows DeepAgent to consolidate its previous thoughts and interaction history into a structured memory schema at any point during its thinking before resuming the agentic reasoning process. This mechanism not only saves tokens and enhances reasoning efficiency over extended interactions but also provides the agent an opportunity to “take a breath”, preventing it from becoming trapped in wrong exploration paths and enabling it to reconsider its strategy. To mitigate information loss during folding, we introduce a brain-inspired memory architecture comprising episodic memory, working memory, and tool memory, all structured with an agent-usable data schema to ensure the stability and utility of the folded memory.
To enhance DeepAgent’s proficiency in mastering these mechanisms, we propose ToolPO, an end-to-end reinforcement learning (RL) training method tailored for general tool use. Existing agentic RL training in general domains presents two significant challenges: (1) The reliance on a multitude of real-world APIs during training can lead to instability, slow execution, and high costs. To prevent this, we leverage LLM-simulated APIs, which enhance the stability and efficiency of the training process. (2) A sparse reward based solely on the final outcome is often insufficient to guarantee the accuracy of intermediate tool calls. We address this by implementing tool-call advantage attribution, which precisely assigns credit to the specific tokens responsible for correct tool invocations, thereby providing a more granular and effective learning signal.
We conduct extensive experiments on a wide range of benchmarks. For (1) General Tool-Use Tasks, we evaluate DeepAgent on ToolBench, API-Bank, TMDB, Spotify, and ToolHop, which feature toolsets scaling from tens to over ten thousand distinct tools. For (2) Downstream Applications, we test its performance on ALFWorld, WebShop, GAIA, and Humanity’s Last Exam (HLE), which require the use of domain-specific toolsets. The overall results in Figure 1 show that DeepAgent achieves superior performance across all scenarios.
Our main contributions are summarized as follows:
  1. (1)
    We propose DeepAgent, the first agentic framework that enables reasoning models to autonomously think, discover tools, and execute actions within a unified reasoning process, empowering LRMs to harness toolsets of arbitrary scale and generalize to complex real-world tasks.
  2. (2)
    We introduce an autonomous memory folding mechanism, complemented by a brain-inspired memory design. This endows the agent with the ability to “take a breath” and reconsider its exploration strategies following unsuccessful attempts.
  3. (3)
    We propose an end-to-end reinforcement learning training methodology for general-purpose tool use, ensuring stability and efficiency in large-scale tool execution during training, as well as accuracy in tool invocation during reasoning.
  4. (4)
    We conduct extensive experiments across eight benchmarks, demonstrating DeepAgent’s superior tool-use capabilities and high adaptability to real-world tasks.
Figure 3. Overview of the DeepAgent framework. The main reasoning model autonomously discovers tools, executes actions, and folds previous memory to restart with structured memories, all within a unified thinking process. The DeepAgent is trained end-to-end with ToolPO, an RL method that uses a tool simulator to simulate large-scale real-world tool APIs, and rewards both final task success and correct intermediate tool calls through fine-grained advantage attribution.
2. Related Work
2.1. Large Reasoning Models
Large Reasoning Models (LRMs) (DeepSeek-AI et al., 2025; Jaech et al., 2024) have demonstrated significant performance improvements in mathematical, scientific, and coding tasks by employing step-by-step slow thinking processes before generating final responses. Existing research has explored various approaches to elicit extended Chain-of-Thought (CoT) reasoning (Wei et al., 2022) from models, including data synthesis for Supervised Fine-Tuning (SFT) (Qin et al., 2024a; Wang et al., 2025b), and end-to-end RL (DeepSeek-AI et al., 2025). Additionally, substantial work has investigated optimization strategies for reasoning models, such as advanced RL training algorithms (Wang et al., 2024b) and improving reasoning efficiency (Yang et al., 2025b). However, models relying solely on parametric knowledge face inherent limitations and cannot interact with the real world. Recent studies have begun exploring tool-augmented reasoning approaches, including Search-o1 (Li et al., 2025a), Search-R1 (Jin et al., 2025a), ToRL (Li et al., 2025d), DeepResearcher (Zheng et al., 2025), and SimpleTIR (Xue et al., 2025). However, these methods typically support only a limited set of research-oriented tools, such as web search, page browsing, and code execution, which constrains their applicability to real-world scenarios that demand access to more diverse tools.
2.2. Autonomous Agents
LLM-powered autonomous agents accomplish real-world tasks by invoking external tools to interact with their environment (Qu et al., 2025a; Wu et al., 2025; Dong et al., 2025b; Liu et al., 2025a; Li et al., 2024; Jin et al., 2025d; Li et al., 2025c; Dong et al., 2025c; Jin et al., 2025b; Tan et al., 2025; Wang et al., 2026; Zhang et al., 2025; Song et al., 2026; Hu et al., 2024, 2025b, 2025a). Current agent methodologies, including ReAct (Yao et al., 2022b), Plan-and-Solve (Wang et al., 2023), Reflextion (Shinn et al., 2024), and CodeAct (Wang et al., 2024a), predominantly follow predefined workflows with fixed execution patterns. This rigid structure limits their ability to fully leverage the autonomous decision-making and deep reasoning capabilities of advanced reasoning models. Recent efforts have investigated training LLMs to autonomously invoke tools through data synthesis and SFT methods (Fang et al., 2025; Xiao et al., 2025) and RL training frameworks (Feng et al., 2025; Wang et al., 2025d; Jiang et al., 2025; Liu et al., 2025b; Kang et al., 2025; Sun et al., 2025; Dong et al., 2025d; Chen et al., 2025b; Dong et al., 2025a). However, most existing methods rely on pre-selected, labeled tools, which limit their applicability to real-world scenarios. Real-world tasks are highly variable and require access to diverse toolsets that cannot be predetermined, aligning with the emerging Model Context Protocol (MCP) (Hou et al., 2025) paradigm. Although some prior work has explored tool retrieval mechanisms (Qin et al., 2024b; Wang et al., 2025a; Shi et al., 2025), most approaches conduct only a single upfront retrieval step and incorporate the retrieved tools, with limited exploration of dynamic tool discovery during task execution. Therefore, we aim to develop a deep reasoning agent capable of dynamically discovering and invoking helpful tools from scalable toolsets to address more generalized real-world tasks.
3. Methodology
3.1. Problem Formulation
We frame the agent’s task as a sequential decision-making process. The agent receives a user-provided question $Q$ and an instruction $I$, and interacts with an environment over a series of steps $t=1,\dots,T$ to accomplish the specified goal. The environment provides access to a collection of tools $\mathcal{T}$ at an arbitrary scale.
At each step $t$, the agent’s state $s_{t}$ consists of the history of all previous actions and their resulting observations, i.e., $s_{t}=(a_{1},o_{1},\dots,a_{t-1},o_{t-1})$. The agent, driven by a policy $\pi$ parameterized by $\theta$, selects an action $a_{t}$ based on the current state, the user question, and the instruction:
(1) $a_{t}\sim\pi_{\theta}(\cdot|s_{t},Q,I).$
An action $a_{t}$ can be one of four types:
  • Internal Thought ($a_{t}^{\text{think}}$): A textual reasoning step generated by the LRM to analyze the problem or plan its next steps. The corresponding observation $o_{t}$ is typically empty.
  • Tool Search ($a_{t}^{\text{search}}$): A natural language query $q_{s}$ to find relevant tools from $\mathcal{T}$. The observation $o_{t}$ is a list of retrieved tools.
  • Tool Call ($a_{t}^{\text{call}}$): The invocation of a specific tool $\tau\in\mathcal{T}$ with a set of arguments. The observation $o_{t}$ is the execution result returned by the tool.
  • Memory Fold ($a_{t}^{\text{fold}}$): A special action to compress the interaction history $s_{t}$ into a structured memory summary. The subsequent state $s_{t+1}$ is then initialized with this compressed memory.
The sequence of states, actions, and observations forms a trajectory $\tau=(s_{1},a_{1},o_{1},\dots,s_{T},a_{T},o_{T})$. The process terminates when the agent completes the task or reaches a maximum step limit. Suppose $R(\tau)$ is a reward function that evaluates the overall success of the trajectory $\tau$, the objective is to learn an optimal policy $\pi_{\theta}^{*}$ that maximizes the expected cumulative reward for a given task:
(2) $\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)].$
3.2. Overview of the DeepAgent Framework
As illustrated in Figure 3, the DeepAgent framework is architected around a main reasoning process, which is supported by several auxiliary mechanisms to ensure robustness and efficiency.
  • Main Reasoning Process: The core of DeepAgent is a powerful large reasoning model that drives the entire task-completion process. In a single stream of thought, the LRM autonomously reasons about the task, dynamically discovers necessary tools, executes actions, and manages its own memory. This unified approach departs from traditional, rigid agent workflows, allowing the LRM to maintain a global perspective on the task.
  • Auxiliary Mechanisms: DeepAgent employs an auxiliary LLM to handle complex interactions with large toolsets and manage long histories. This background model enhances system stability by: (1) filtering and summarizing retrieved tool documentation if it’s too lengthy, (2) denoising and condensing verbose information returned from tool calls, and (3) compressing long interaction histories into a structured memory. This division of labor allows the main LRM to concentrate on high-level strategic reasoning.
3.3. Autonomous Tool Search and Calling
DeepAgent’s main LRM performs all actions by generating specific textual prompts within its continuous reasoning process. These actions are then intercepted and executed by the system.
Tool Search
When the agent determines it needs a tool, it generates a tool search query $q_{s}$ encapsulated within special tokens: <tool_search> $q_{s}$ </tool_search>. The system’s tool retriever operates via dense retrieval. First, we build an index by pre-computing an embedding $E(d_{i})$ for the documentation $d_{i}$ of each tool $\tau_{i}\in\mathcal{T}$ using an embedding model $E$. During inference, given the query $q_{s}$, the system retrieves the top-$k$ tools by ranking them based on the cosine similarity $\text{sim}(\cdot,\cdot)$:
(3) $\mathcal{T}_{\text{retrieved}}=\underset{\tau_{i}\in\mathcal{T}}{\text{top-k}}\left(\text{sim}\left(E\left(q_{s}\right),E\left(d_{i}\right)\right)\right).$
The retrieved tool documentation is then processed by the auxiliary LLM —summarized if too lengthy, otherwise provided directly—and returned to the main LRM’s context: <tool_search_result> relevant tools </tool_search_result>.
Tool Call
To execute a tool, the agent generates a structured call including the tool’s name and arguments: <tool_call> {”name”: ”tool_name”, ”arguments”: …} </tool_call>. The framework parses this call, executes the tool, and captures the output. This output is, if necessary, summarized by the auxiliary LLM to ensure it is concise and helpful, before being fed back into the reasoning context: <tool_call_result> helpful information </tool_call_result>.
3.4. Autonomous Memory Folding and Brain-Inspired Memory Schema
The agent can trigger memory folding at any logical point in its reasoning process—such as after completing a sub-task or realizing an exploration path was incorrect—by generating a special token: <fold_thought>. Upon detecting this token, the system initiates the memory folding process. The auxiliary LLM (parameterized by $\theta_{\text{aux}}$) processes the entire preceding interaction history $s_{t}$ and generates three structured memory components in parallel:
(4) $(M_{E},M_{W},M_{T})=f_{\text{compress}}(s_{t};\theta_{\text{aux}}).$
These compressed episodic ($M_{E}$), working ($M_{W}$), and tool ($M_{T}$) memories then replace the raw interaction history, enabling the agent to proceed with a refreshed and condensed view of its progress while avoiding entrapment in incorrect exploration paths.
Inspired by human cognitive systems, the structured memory $M_{t}$ is composed of three distinct components that are generated in parallel: $M_{t}=(M_{E},M_{W},M_{T})$, where $M_{E},M_{W},M_{T}$ denote episodic, working, and tool memories, respectively.
  • Episodic Memory ($M_{E}$): This component serves as a high-level log of the task, recording key events, major decision points, and sub-task completions. It provides the agent with long-term context regarding the overall task structure and its overarching goals.
  • Working Memory ($M_{W}$): This contains the most recent information, such as the current sub-goal, obstacles encountered, and near-term plans. It is the core component that ensures the continuity of the agent’s reasoning across the memory fold.
  • Tool Memory ($M_{T}$): This consolidates all tool-related interactions, including which tools have been used, how they were invoked, and their effectiveness. It allows the agent to learn from its experiences, refining its tool selection and usage strategies.
To ensure that the compressed memory is stable and easily parsed by the agent, we employ an agent-usable data schema in JSON format instead of unstructured natural language. It offers two main benefits: maintaining a controllable and predictable structure, and mitigating the loss of critical details that can occur when summarizing long-form text. Details of the data schema are in Appendix C.
Table 1. Main results on general tool usage tasks, encompassing scenarios with both labeled tools and open-set tool retrieval over large-scale toolsets. We report Pass@1 metric for all tasks. For 32B models, the best results are in bold and the second are underlined. Results from larger or closed-sourced models are in gray color for reference.
Method Backbone ToolBench API-Bank TMDB Spotify ToolHop
Success Path Success Path Success Path Success Path Correct Path
Scenario 1: Completing Tasks w/ Ground-truth Tools
Workflow-based Methods
ReAct Qwen2.5-32B 41.0 64.7 60.4 68.3 46.0 65.3 29.8 56.3 37.6 49.1
CodeAct Qwen2.5-32B 53.0 68.3 62.4 70.6 48.0 67.4 33.3 58.7 34.7 48.8
Plan-and-Solve Qwen2.5-32B 52.0 65.4 58.4 67.5 51.0 71.6 28.1 54.8 39.2 49.7
ReAct QwQ-32B 52.0 61.6 73.3 78.6 43.0 65.3 47.4 69.4 47.4 51.6
CodeAct QwQ-32B 54.0 63.4 74.3 79.4 55.0 74.5 52.6 75.4 43.2 53.4
Plan-and-Solve QwQ-32B 55.0 64.7 70.3 75.4 48.0 61.3 49.1 70.6 45.4 50.6
ReAct Qwen2.5-72B 56.0 69.3 73.3 78.6 47.0 67.7 57.9 76.6 44.8 55.4
ReAct GPT-4o 52.0 53.9 79.2 83.3 77.0 89.3 47.4 70.6 40.0 53.7
ReAct DeepSeek-R1 57.0 68.3 71.3 76.2 76.0 89.0 64.9 81.3 50.2 61.8
Autonomous Tool Usage within Reasoning
DeepAgent-32B-Base QwQ-32B 63.0 74.3 76.2 81.0 85.0 92.0 70.2 89.3 49.1 59.8
DeepAgent-32B-RL QwQ-32B 69.0 78.6 75.3 80.2 89.0 94.8 75.4 92.0 51.3 62.5
Scenario 2: Completing Tasks w/ Open-Set Tool Retrieval
Workflow-based Methods
ReAct Qwen2.5-32B 55.0 20.8 16.0 42.0 11.0 34.5 7.0 25.4 13.2 17.9
CodeAct Qwen2.5-32B 51.0 19.0 22.0 49.6 19.0 46.8 10.5 31.6 12.7 17.4
Plan-and-Solve Qwen2.5-32B 54.0 20.4 18.0 42.8 15.0 40.5 8.8 26.3 12.0 16.3
ReAct QwQ-32B 44.0 19.0 20.0 52.7 18.0 40.3 22.8 45.5 27.1 22.3
CodeAct QwQ-32B 48.0 21.6 16.0 45.0 31.0 52.8 24.6 49.6 29.0 26.1
Plan-and-Solve QwQ-32B 45.0 19.6 18.0 44.3 24.0 46.8 19.3 42.7 25.7 20.8
ReAct Qwen2.5-72B 52.0 21.6 14.0 38.9 28.0 50.7 21.1 48.5 21.1 19.9
ReAct GPT-4o 41.0 28.9 18.0 42.8 35.0 56.8 17.5 26.3 24.1 28.6
ReAct DeepSeek-R1 47.0 22.3 12.0 57.3 34.0 53.1 29.8 51.7 36.2 32.9
Autonomous Tool Retrieval and Usage within Reasoning
DeepAgent-32B-Base QwQ-32B 60.0 35.7 22.0 61.8 52.0 71.8 49.1 68.6 38.4 40.3
DeepAgent-32B-RL QwQ-32B 64.0 37.2 24.0 64.9 55.0 74.3 50.9 74.4 40.6 40.5
Table 2. Main results on downstream task applications, spanning Embodied AI (ALFWorld), Online Shopping (WebShop), General AI Assistants (GAIA), and Humanity’s Last Exam (HLE). We report Pass@1 for all tasks. For 32B models, the best results are in bold and the second are underlined. Results from larger or closed-sourced models are in gray color for reference.
Method Backbone ALFWorld WebShop GAIA HLE
Success Path Success Score Text $M$ File All Text $M$ All
Completing Tasks w/ Task-specific Toolsets
Workflow-based Methods
ReAct Qwen2.5-32B 60.4 79.1 6.0 28.8 25.2 16.7 13.2 21.2 6.5 7.1 6.6
CodeAct Qwen2.5-32B 65.7 83.3 12.4 34.5 28.2 20.8 18.4 24.8 7.5 8.0 7.6
Reflextion Qwen2.5-32B 66.4 86.0 9.2 31.6 29.1 20.8 18.4 25.5 5.9 5.3 5.8
Plan-and-Solve Qwen2.5-32B 63.4 80.4 7.6 29.3 27.2 16.7 15.8 23.0 7.2 6.2 7.0
ReAct QwQ-32B 82.1 87.8 17.2 45.3 35.0 8.3 36.8 31.5 13.2 8.8 12.2
CodeAct QwQ-32B 78.4 86.2 18.0 46.4 38.8 20.8 31.6 34.5 14.2 8.0 12.8
Reflextion QwQ-32B 85.1 88.4 21.6 50.4 37.9 20.8 36.8 35.2 11.9 7.1 10.8
Plan-and-Solve QwQ-32B 79.1 84.7 16.0 43.8 36.9 16.7 34.2 33.3 12.9 9.7 12.2
AgentLM* Llama2-70B 86.0 - - 64.9 - - - - - - -
ReAct Qwen2.5-72B 86.5 86.5 22.0 44.5 32.0 20.8 31.6 30.3 9.0 8.0 8.8
ReAct DeepSeek-R1 79.1 85.8 19.6 49.7 43.7 29.2 39.5 40.6 14.2 8.8 13.0
ReAct GPT-4o 65.7 87.8 15.6 52.5 35.0 16.7 36.8 32.7 13.2 10.6 12.6
ReAct Claude-4 93.3 91.5 20.4 56.6 56.3 37.5 52.6 52.7 15.5 16.8 15.8
Autonomous Tool Usage within Reasoning
Deep Research OpenAI (o3) - - - - - - - 67.4 - - 26.6
WebThinker QwQ-32B - - - - 48.5 25.0 13.2 37.0 14.2 8.8 13.0
HiRA QwQ-32B 84.3 87.6 23.2 51.9 44.7 33.3 42.1 42.5 14.5 10.6 13.6
DeepAgent-32B-Base QwQ-32B 88.1 91.4 32.0 55.4 49.5 37.5 44.7 46.7 19.1 13.3 17.8
DeepAgent-32B-RL QwQ-32B 91.8 92.0 34.4 56.3 58.3 33.3 52.6 53.3 21.7 15.0 20.2
3.5. End-to-end RL Training with ToolPO
We train DeepAgent end-to-end with Tool Policy Optimization (ToolPO), an RL approach designed for general tool-using agents.
Training Data Collection
We first collect a diverse training dataset spanning four categories. To instill general tool-use capabilities, we use ToolBench (Qin et al., 2024b). For real-world interaction, we leverage ALFWorld (Shridhar et al., 2021) and WebShop (Yao et al., 2022a). To enhance deep research skills, we incorporate data from WebDancer (Wu et al., 2025) and WebShaperQA (Tao et al., 2025). Lastly, to improve mathematical reasoning with code, we use DeepMath (He et al., 2025). Further details are available in Appendix A.1.
Tool Simulator
Training an agent that interacts with thousands of real-world APIs is often impractical due to instability, latency, and cost. To address this, we develop an LLM-based Tool Simulator. This simulator, powered by an auxiliary LLM, mimics the responses of real-world APIs (e.g., RapidAPI). This approach provides a stable, efficient, and low-cost environment for robust RL training.
Global and Tool-Call Advantage Attribution
For each input prompt, we sample a group of $K$ trajectories $\{\tau_{1},\dots,\tau_{K}\}$. ToolPO defines two distinct reward components. The first is a reward for overall task success, $R_{\text{succ}}(\tau)$, which is a task-success score that reflects the quality of the final outcome (e.g., the accuracy of the final answer). The second is a tool-call reward, $R_{\text{action}}(\tau)$, which reflects the quality of intermediate actions. This action-level reward is composed of rewards for correct tool invocations and efficient memory folding. Specifically, $R_{\text{action}}(\tau)=\lambda_{1}\sum_{t=1}^{T}C(a_{t}^{\text{call}})+\lambda_{2}S_{\text{pref}}(\tau)$, where $C(a_{t}^{\text{call}})$ is 1 if a tool call is correct and 0 otherwise. $S_{\text{pref}}(\tau)$ is a preference score encouraging efficient use of memory folding, defined by comparing a trajectory with folding ($\tau_{\text{fold}}$) to one without ($\tau_{\text{direct}}$): $S_{\text{pref}}=(L(\tau_{\text{direct}})-L(\tau_{\text{fold}}))/(L(\tau_{\text{direct}})+L(\tau_{\text{fold}}))$.
Based on these rewards, we compute two separate group-relative advantages. The task success advantage for trajectory $\tau_{k}$ is:
(5) $A_{\text{succ}}(\tau_{k})=R_{\text{succ}}(\tau_{k})-\frac{1}{K}\sum\nolimits_{j=1}^{K}R_{\text{succ}}(\tau_{j}),$
which is attributed to all generated tokens in $\tau_{k}$, providing a global learning signal. Similarly, the action-level advantage is:
(6) $A_{\text{action}}(\tau_{k})=R_{\text{action}}(\tau_{k})-\frac{1}{K}\sum\nolimits_{j=1}^{K}R_{\text{action}}(\tau_{j}).$
Crucially, this advantage is attributed only to the specific tokens that constitute the tool call and memory folding actions. This fine-grained credit assignment provides a more targeted signal for learning correct and efficient tool use.
Optimization Objective
The total advantage for a given token $y_{i}$ in trajectory $\tau_{k}$ is the sum of the global and local advantages:
(7) $A(y_{i})=A_{\text{succ}}(\tau_{k})+M(y_{i})\cdot A_{\text{action}}(\tau_{k}),$
where $M(y_{i})$ is a mask that is 1 if $y_{i}$ is part of a tool-call or memory-fold token sequence, and 0 otherwise. ToolPO then optimizes the policy using a clipped surrogate objective function:
(8) $\mathcal{L}_{\text{ToolPO}}(\theta)=$
$\mathbb{E}_{\tau_{k}}\left[\sum\nolimits_{i=1}^{|\tau_{k}|}\min\Big(\rho_{i}(\theta)A(y_{i}),\text{clip}(\rho_{i}(\theta),1-\epsilon,1+\epsilon)A(y_{i})\Big)\right].$
Here, $\rho_{i}(\theta)=\frac{\pi_{\theta}(y_{i}|y_{<i},s)}{\pi_{\theta_{\text{old}}}(y_{i}|y_{<i},s)}$ is the probability ratio for token $y_{i}$. This objective encourages the model to increase the probability of both intermediate actions and end-to-end task accomplishment that exhibit positive relative advantage, thereby ensuring stable and effective policy updates.
4. Experimental Settings
4.1. Tasks and Datasets
We conduct extensive experiments on a wide range of benchmarks, including general tool-use and downstream applications.
General Tool-Use. These benchmarks encompass a broad range of distinct tools, scaling from tens to over ten thousand, making them ideal for evaluating the scalability of different approaches. We utilize four representative scenarios: ToolBench (Qin et al., 2024b), based on over 16,000 real-world APIs, for which we use the G3 subset requiring multi-step, multi-tool calls; API-Bank (Li et al., 2023), which includes 314 human-annotated dialogues with 73 APIs and 753 API calls, to assess planning, retrieval, and calling capabilities; RestBench (Song et al., 2023), comprising scenarios from the TMDB movie database (54 tools, avg. 2.3 calls/question) and the Spotify music player (40 tools, avg. 2.6 calls/question) to simulate typical REST applications; and ToolHop (Ye et al., 2025), a multi-hop reasoning dataset with 3,912 locally executable tools that necessitate 3 to 7 sequential tool calls per task. For these tasks, we adopt two settings: given ground-truth tools and given entire toolsets with tool retrieval capabilities.
Downstream Applications. We also evaluate downstream applications with domain-specific toolsets: ALFWorld (Shridhar et al., 2021) (text embodied tasks with nine actions, e.g., move/take), WebShop (Yao et al., 2022a) (shopping with ‘search’ and ‘click’), GAIA (Mialon et al., 2024) (web search/browsing, VQA, code, file reading), and Humanity’s Last Exam (HLE) (Phan et al., 2025) (code, search, browsing, VQA). These tasks test long-horizon interaction in more realistic environments, requiring state tracking, error recovery, and coordination across heterogeneous tools; we equip agents with task-specific toolsets.
4.2. Baselines
Our baselines include: (1) Workflow-based Methods: ReAct (Yao et al., 2022b) alternates explicit reasoning with environment actions in a Reason-Act-Observe loop. CodeAct (Wang et al., 2024a) expresses actions as executable Python code that runs in an interpreter. Plan-and-Solve (Wang et al., 2023) first sketches a high-level plan and then executes it step by step. Reflexion (Shinn et al., 2023) enhances learning through verbal self-reflection after failed attempts. AgentLM (Zeng et al., 2024) uses instruction tuning to enhance general agent capabilities of LLMs. (2) Autonomous Tool Usage within Reasoning: WebThinker (Li et al., 2025b) interleaves thinking with web search and deep web exploration. HiRA (Jin et al., 2025c) introduces a hierarchical agent architecture where a meta planner decomposes tasks, a coordinator routes subtasks, and specialized executors solve them with dual-channel memory. OpenAI Deep Research (OpenAI, 2025) is an agentic system based on reasoning models.
4.3. Implementation Details
We use QwQ-32B (Team, 2024) as DeepAgent’s backbone model, with Qwen2.5-32B-Instruct (Qwen et al., 2024) as the auxiliary model in our main results. Text generation employs a maximum of 81,920 tokens with temperature 0.7, top_p 0.8, top_k 20, and repetition penalty 1.05. Web search and page browsing are implemented using Google Serper API and Jina Reader API, respectively. The VQA tool is based on Qwen2.5-VL-32B-Instruct (Bai et al., 2025). Tool retrieval is performed using bge-large-en-v1.5 (Xiao et al., 2024). Training consists of 100 steps of ToolPO with batch size 64, $\lambda_{1}=\lambda_{2}=1$, rollout size $K=8$, and maximum sequence length 32,768. Additional details are provided in Appendix B. All experiments are conducted on 64 NVIDIA H20-141GB GPUs.
Figure 4. Visualization of training dynamics, including (a) reward scores and (b) validation scores across training steps.
Table 3. Ablation studies on the components of DeepAgent, where the best results are in bold.
Method Tool-Usage Application Avg.
ToolB. ToolH. WebS. GAIA
DeepAgent-32B-RL 64.0 40.6 34.4 53.3 48.1
w/o Training (Base) 60.0 38.4 32.0 46.7 44.3
w/o Memory Folding 63.0 36.6 32.4 44.7 44.2
w/o Tool Simulation 62.0 35.2 33.6 48.5 44.8
w/o Tool Adv. Attribution 62.0 39.6 33.2 49.5 46.1
5. Experimental Results
5.1. Main Results on General Tool Usage Tasks
Table 1 presents the results on general tool usage, leading to several key observations. (1) DeepAgent’s End-to-End Reasoning Surpasses Workflow-Based Methods. DeepAgent’s holistic agentic process consistently outperforms rigid, predefined workflows. For instance, on labeled-tool tasks, DeepAgent-32B-RL achieves success rates of 89.0% on TMDB and 75.4% on Spotify, substantially exceeding the strongest 32B baseline scores of 55.0% and 52.6%, respectively. This underscores the benefit of a holistic agentic process over rigid, predefined action cycles. (2) DeepAgent Maintains Robustness in Open-Set Scenarios. This advantage is more pronounced in open-set scenarios where dynamic tool discovery is critical. On ToolBench and ToolHop, DeepAgent-32B-RL achieves success rates of 64.0% and 40.6%, respectively, far exceeding the top baseline scores of 54.0% and 29.0%. This demonstrates that DeepAgent’s strategy of dynamically discovering tools as needed within the reasoning process is far more robust and scalable in realistic open-set scenarios. (3) ToolPO Training Further Improves Tool-Usage Capabilities. The proposed ToolPO RL strategy provides significant further gains. The trained DeepAgent-32B-RL model consistently improves upon its base version, boosting success rates on ToolBench by up to 6.0% and on Spotify (labeled) by 5.2%. This validates the effectiveness of the ToolPO strategy, which uses an LLM-based tool simulator and fine-grained advantage attribution.
5.2. Main Results on Downstream Applications
Table 2 shows the results on downstream applications, which require agents to handle long-horizon interactions in complex environments. (1) The autonomous reasoning paradigm generally outperforms workflow-based methods. On complex application tasks, methods that integrate tool usage into continuous reasoning consistently outperform rigid, predefined workflows. On GAIA, both DeepAgent-32B-Base (46.7) and HiRA (42.5) significantly exceed the best workflow-based method CodeAct (34.5). Similarly, on WebShop, DeepAgent-32B-Base (32.0) substantially surpasses CodeAct (18.0). This demonstrates that long-horizon interaction tasks require deep agentic reasoning capabilities to achieve superior task accomplishments. (2) DeepAgent demonstrates superior performance across various application tasks. DeepAgent achieves state-of-the-art performance among 32B models. On GAIA, DeepAgent-32B-RL scores 53.3 vs. HiRA’s 42.5, and on ALFWorld reaches 91.8% vs. HiRA’s 84.3%. This stems from DeepAgent’s seamless integration of actions into coherent reasoning, enabling end-to-end execution with autonomous memory folding, which is advantages unavailable to workflow-constrained methods. (3) ToolPO training further improves performance on downstream applications. ToolPO training yields consistent gains over the base model. DeepAgent-32B-RL improves GAIA scores from 46.7 to 53.3 (+6.6) and ALFWorld success rates from 88.1% to 91.8% (+3.7), demonstrating that ToolPO effectively enhances reasoning and tool usage capabilities for complex task completion.
Table 4. Effectiveness analysis of autonomous tool retrieval strategy in open-set scenarios compared to pre-retrieved tool methods. Numbers in parentheses indicate toolset sizes.
Method ToolB. ToolH. TMDB Spotify Avg.
(16k) (3.9k) (54) (40)
ReAct Workflow
Input Retrieved Tool 35.0 25.4 14.0 15.0 22.4
Auto. Tool Retrieval 34.0 37.1 18.0 27.8 28.0
Plan-and-Solve Workflow
Input Retrieved Tool 37.0 24.8 19.0 16.0 24.2
Auto. Tool Retrieval 45.0 25.7 24.0 19.3 28.5
End-to-end Agentic Reasoning (DeepAgent)
Input Retrieved Tool 53.0 37.0 34.0 43.9 42.0
Auto. Tool Retrieval 64.0 40.6 55.0 50.9 52.6
5.3. Analysis of Training Dynamics
Figure 4 shows the training dynamics of DeepAgent, including the reward scores and validation scores across training steps. As shown in the figure, (1) DeepAgent trained with ToolPO achieves higher upper bounds on both reward and validation scores compared to the commonly used GRPO. (2) Moreover, the training reward exhibits less fluctuation than GRPO, demonstrating better training stability. This indicates that using tool simulators instead of directly training with unstable real-world APIs, along with employing tool-call process supervision, enables more stable and effective training of tool-usage capabilities.
Figure 5. Scaling analysis of performance with respect to maximum action limits on WebShop and GAIA datasets.
5.4. Ablation Studies
We conduct ablation studies in Table 3 to validate the effectiveness of each component in DeepAgent. (1) Importance of ToolPO Training: Removing ToolPO training (the Base model) results in the most significant performance drop (from 48.1 to 44.3). This highlights the central role of our end-to-end RL method in enhancing tool use and complex task completion. (2) Effectiveness of Memory Folding: The absence of memory folding also leads to a substantial performance decline (average score drops to 44.2), particularly on the long-horizon task GAIA (from 53.3 to 44.7). This confirms that the autonomous memory folding mechanism, allowing the agent to ”take a breath” and replan, is crucial for robust long-term interaction. (3) Contribution of Training Strategies: Removing the tool simulator and tool-call advantage attribution both lead to performance degradation. This validates that the tool simulator enables more stable training, and fine-grained advantage attribution provides precise learning signals.
5.5. Effectiveness of Tool Retrieval Strategies
To compare pre-retrieving tools versus autonomous discovery during task execution, we conduct experiments shown in Table 4. We find: (1) The on-demand nature of dynamic tool discovery yields superior performance and robust scalability. Autonomous tool retrieval during reasoning consistently outperforms pre-retrieved tools across all frameworks, demonstrating the superiority of on-demand tool access in open-set scenarios. Performance gains are most pronounced on large toolsets like ToolBench (16k tools) and ToolHop (3.9k tools), indicating robust scalability for real-world tasks. (2) DeepAgent synergizes better with dynamic retrieval. Combined with autonomous tool retrieval, our framework achieves the best results by a large margin, scoring 52.6 on average versus 28.5 for the best workflow-based method. This demonstrates that DeepAgent’s architecture is uniquely suited for dynamic tool discovery.
5.6. Scaling Analysis of Action Limits
Figure 5 illustrates the performance of DeepAgent and ReAct on the WebShop and GAIA datasets as the maximum action limit is varied. The results yield several key insights. (1) DeepAgent consistently and significantly outperforms the ReAct baseline across all tested action limits on both datasets, demonstrating its superior effectiveness. (2) For both agents, performance generally improves as the maximum number of actions increases. This suggests that complex tasks benefit from a longer interaction horizon, allowing for more thorough exploration and reasoning. (3) DeepAgent exhibits stronger scalability. As the action limit increases, the performance gap between DeepAgent and ReAct widens, particularly on WebShop. This sustained gain suggests DeepAgent strategically selects effective, task-relevant actions, avoiding the wasteful steps that limit ReAct’s scalability.
Table 5. Performance with different reasoning model backbones: MOE-based models with 30B and 235B parameters.
Method Tool-Usage Application Avg.
ToolB. ToolH. ALF. WebS. GAIA
Qwen3-30B-A3B-Thinking
ReAct 52.0 22.0 67.9 18.4 34.5 35.7
Plan-and-Solve 50.0 23.6 68.7 20.4 35.2 37.0
DeepAgent (Base) 59.0 47.5 69.4 31.4 39.4 46.9
Qwen3-235B-A22B-Thinking
ReAct 61.0 40.9 79.9 21.6 36.4 45.1
Plan-and-Solve 63.0 43.0 78.4 24.4 38.4 46.0
DeepAgent (Base) 67.0 48.2 85.8 37.2 51.5 55.7
5.7. Generalization Across Different Backbones
Table 5 shows the performance of DeepAgent with different backbone large reasoning models, including Qwen3-30B-A3B-Thinking and Qwen3-235B-A22B-Thinking (Yang et al., 2025a). (1) DeepAgent consistently outperforms workflow-based methods. With both the 30B and 235B MoE-based reasoning models as backbones, DeepAgent maintains a significant performance margin over ReAct and Plan-and-Solve, demonstrating the generalizability of its agentic reasoning approach. (2) DeepAgent scales effectively with larger models. While all methods benefit from scaling the backbone from a 30B to a 235B model, DeepAgent shows the largest absolute performance gains on complex application tasks.
6. Conclusion
In this work, we introduce DeepAgent, an end-to-end reasoning agent that unifies thinking, tool discovery, and execution into a single, coherent agentic reasoning process. To enable robust long-horizon interaction, we propose an autonomous memory folding mechanism that compresses interaction history into a structured memory, allowing the agent to ”take a breath” and reconsider its strategy. We also introduce ToolPO, an end-to-end RL method that leverages LLM simulated APIs for stable training and fine-grained advantage attribution for precise credit assignment to tool invocations. Extensive experiments on general tool-use and downstream applications demonstrate that DeepAgent significantly outperforms various baseline agents, particularly in open-set scenarios requiring dynamic tool discovery over scalable toolsets. This work opens new avenues for creating more general and capable agents, with future research directions including developing advanced memory management, multimodal (or even omni-modal) capabilities, and human-agent collaborative reasoning.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China No. 62272467, and the China Postdoctoral Science Foundation under Grant Number 2025T180440. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
Appendix
Appendix A Datasets
A.1. Training Data
We collected a diverse training dataset spanning four task categories to instill comprehensive agent capabilities.
  • General Tool-Use: We sample 1k instances for labeled-tool scenarios and 1k for tool-retrieval from the ToolBench (Qin et al., 2024b) training set. This data is intended to instill a generalized ability to use diverse tools and leverage large toolsets through retrieval.
  • Real-World Interaction: We utilize 500 instances from ALFWorld (Shridhar et al., 2021) and 500 from WebShop (Yao et al., 2022a), sampled from their training sets, to teach the model to interact effectively with environments, manage state transitions, and achieve user goals.
  • Deep Research: We include 200 instances from WebDancer (Wu et al., 2025) and 500 from WebShaperQA (Tao et al., 2025) to enhance the model’s proficiency in using web search and page browsing for in-depth information gathering.
  • Mathematical Reasoning: We collect 0.9k problems from the DeepMath dataset (He et al., 2025) to strengthen the model’s ability to use code as a tool for complex mathematical computations.
A.2. Benchmarks
We conduct extensive experiments on a wide range of benchmarks, including general tool-use and downstream applications.
General Tool-Use
These benchmarks encompass a broad range of distinct tools (from tens to over ten thousand), thus offering a testbed for evaluating different approaches to toolset scaling.
  • ToolBench (Qin et al., 2024b): A large-scale benchmark containing over 16,000 real-world REST APIs spanning 49 categories. Test subsets include 100 test cases, designed to evaluate LLMs in both single-tool and complex multi-tool scenarios.
  • API-Bank (Li et al., 2023): A comprehensive benchmark for tool-augmented LLMs. It features a runnable evaluation system with 73 API tools and a large training set (over 2,200 dialogues across 2,211 APIs from 1,008 domains), assessing LLMs’ capabilities in planning, retrieving, and calling APIs.
  • TMDB (Song et al., 2023): A sub-scenario of RestBench focused on the TMDB movie database, consisting of 100 questions that utilize 54 local tools and require an average of 2.3 sequential API calls.
  • Spotify (Song et al., 2023): A sub-scenario of RestBench simulating a Spotify music player, featuring 57 questions and 40 local tools, demanding an average of 2.6 sequential API calls to complete the tasks.
  • ToolHop (Ye et al., 2025): A multi-hop reasoning dataset comprising 995 complex questions. It leverages 3,912 locally executable tools and requires between 3 to 7 sequential tool calls per task.
Downstream Applications
These benchmarks test the capability of different approaches in handling complex real-world tasks, which often require the use of domain-specific toolsets.
  • ALFWorld (Shridhar et al., 2021): A benchmark for simple Embodied AI tasks set in a text environment. Agents must complete objectives using a finite set of low-level embodied actions (eg., move, take) to test navigation and object manipulation.
  • WebShop (Yao et al., 2022a): A challenging online shopping environment that provides 12,087 crowd-sourced tasks over a catalog of 1.18 million products. Agents interact with the simulated e-commerce website using core APIs: search[Query] and choose[Text Button].
  • GAIA (Mialon et al., 2024): A complex benchmark for General AI Assistants, consisting of 466 real-world questions (with a 300-question held-out test set). It requires the flexible application of a broad general-purpose toolset including web browsing, code execution, multi-modal processing, and file handling.
  • Humanity’s Last Exam (HLE) (Phan et al., 2025): A benchmark featuring 2,500 highly difficult, multi-disciplinary questions (graduate-level). It primarily evaluates the model’s intrinsic deep reasoning and multi-modal understanding capabilities, as the questions are designed to be insoluble by simple external search tools.
Appendix B Implementation Details
For DeepAgent, we use QwQ-32B (Team, 2024) as the main reasoning model for the results in Table 1 and Table 2, and Qwen3-30B-A3B-Thinking-2507 (Yang et al., 2025a) with Qwen3-235B-A22B-Thinking-2507 (Yang et al., 2025a) in Table 5. We use Qwen2.5-32B-Instruct (Qwen et al., 2024) as the auxiliary model for (1) filtering lengthy tool search results and execution outputs (this is also applied to all baselines), (2) simulating RapidAPIs during ToolPO training, and (3) generating folded memory from interaction history. For the baselines, we use either QwQ-32B or Qwen2.5-32B-Instruct as the backbone model. Text generation for all models uses a maximum of 81,920 tokens, with a temperature of 0.7, top_p of 0.8, top_k of 20, and a repetition penalty of 1.05. The maximum number of actions is set to 50.
Web search and page browsing are implemented using the Google Serper API and Jina Reader API, respectively. The VQA tool is based on Qwen2.5-VL-32B-Instruct (Bai et al., 2025), which takes a question and an image as input and outputs a model-generated response. Tool retrieval is performed using bge-large-en-v1.5 (Xiao et al., 2024). All tool documentation follows the standard OpenAI function definition format: {”name”: ”…”, ”description”: ”…”, ”parameters”: {”type”: ”object”, ”properties”: {”param1”: {”type”: ”…”, ”description”: ”…”}, …, ”required”: [”param1”]}}. This format is used for building the toolset index and for all prompts given to the agents.
Training consists of 100 steps of ToolPO with a batch size of 64, $\lambda_{1}=\lambda_{2}=1$, rollout size $K=8$, and a maximum sequence length of 32,768. The maximum number of actions is 50. The training framework is based on VeRL (Sheng et al., 2024) for multi-node distributed training. All experiments are conducted on 64 NVIDIA H20-141GB GPUs.
Appendix C Memory Schema
Our brain-inspired memory architecture contains three components: episodic, working, and tool memory. To support stable memory folding and reduce information loss, we define each component with a fixed JSON schema, enabling reliable parsing and use of compressed memories for long-horizon reasoning.
Episodic Memory Schema
Episodic memory records high-level task progression (milestones, decisions, outcomes) to preserve long-term context. The format is: {”task_description”: ”A general summary of what the reasoning history has been doing and the overall goals it has been striving for.”, ”key_events”: [{”step”: ”step number”, ”description”: ”A detailed description of the specific action taken, decision made, or milestone achieved at this step, including relevant context and reasoning behind the choice.”, ”outcome”: ”A detailed account of the direct result, observation, or feedback received from this action or decision, including any new information gained or changes in the task state.”}], ”current_progress”: ”A general summary of the current progress of the task, including what has been completed and what is left to be done.”}
Working Memory Schema
Working memory captures the immediate goal, active challenges, and next actions to maintain continuity across folds. The format is: {”immediate_goal”: ”A clear summary of the current subgoal—what you are actively working toward at this moment.”, ”current_challenges”: ”A concise summary of the main obstacles or difficulties you are presently encountering.”, ”next_actions”: [{ẗype”: ”tool_call or planning ordecision”, d̈escription”: ”Anticipate and describe the next concrete action you intend to take to advance the task.”}]}
Tool Memory Schema
Tool memory consolidates tool-use experience (success rates, effective parameters, common errors) and derives reusable rules. The format is: {”tools_used”: [{”tool_name”: ”string”, ”success_rate”: ”float”, ”effective_parameters”: [”param1”, ”param2”], ”common_errors”: [”error_type1”, ”error_type2”], ”response_pattern”: ”description of typical output”, ”experience”: ”Reflect and summarize your experience, including both successes and failures.”}], ”derived_rules”: [”When X condition occurs, prefer tool Y”, ”Tool Z works best with parameter A set to B”]}
Appendix D Case Study
To illustrate the effectiveness of our DeepAgent framework in handling complex, multi-step tasks that require coordinated use of multiple tools, we present a detailed case in Table 6. This example demonstrates how DeepAgent autonomously navigates tool selection, executes sequential actions, and synthesizes results to provide comprehensive solutions to user queries.
Table 6. A case study of DeepAgent for multi-step, multi-tool tasks on ToolBench, showing only key actions.
Example of DeepAgent on a Multi-Tool Task
User Query:
I’m organizing a film festival and I need assistance in finding the best films. Can you search for videos related to ’documentary’ on Vimeo? Additionally, fetch the related people in the ’cinema’ category to invite them as guest speakers. Finally, provide me with a streaming link for a YouTube video with the ID ’UxxajLWwzqY’.
DeepAgent Processing Output
Okay, I need to solve three tasks: find documentaries on Vimeo, find cinema-related people on Vimeo, and get a YouTube streaming link. I’ll start by searching for the right tools.
¡tool_search¿ Vimeo video search for documentaries ¡/tool_search¿ The search returned several tools. search_videos and get_videos_by_tag seem most relevant. I’ll use the search tool first to find documentaries.
¡tool_call¿ {”name”: ”search_videos”, ”arguments”: {”query”: ”documentary”, ”sort”: ”most_liked”}} ¡/tool_call¿ The tool returned a list of popular documentary videos. Now for the second task: finding people in the ’cinema’ category. Since there is no direct tool to search for people, I will try searching for videos with the ’cinema’ tag and identify their creators.
¡tool_call¿ {”name”: ”get_videos_by_tag”, ”arguments”: {”tag”: ”cinema”, ”sort”: ”most_liked”}} ¡/tool_call¿ This call returned videos tagged with ’cinema’ and included user information for each video, such as ”John Doe” and ”Jane Smith”, who are potential guest speakers.
Finally, for the YouTube streaming link, I’ll first verify the video ID exists using the YouTube search tool.
¡tool_call¿ {”name”: ”search_for_youtube_v3”, ”arguments”: {”part”: ”snippet”, ”type”: ”video”, ”q”: ”UxxajLWwzqY”}} ¡/tool_call¿ The tool confirmed the video ID is valid. I can now construct the streaming link. I will now consolidate all the information to answer the user’s query.
Final Answer:
1. Documentaries on Vimeo: Found nature and space documentaries.
2. Cinema Guest Speakers: John Doe, Jane Smith with Vimeo profiles.
3. YouTube Link: https://www.youtube.com/watch?v=UxxajLWwzqY